### v1

# UniRVQA: A Unified Framework for Retrieval-Augmented Vision Question Answering via Self-Reflective Joint Training

Abstract.
Knowledge-based Vision Question Answering (KB-VQA) systems address complex visual-grounded questions requiring external knowledge, such as web-sourced encyclopedia articles. Existing methods often use sequential and separate frameworks for the retriever and the generator with limited parametric knowledge sharing. However, since both retrieval and generation tasks require accurate understanding of contextual and external information, such separation can potentially lead to suboptimal system performance. Another key challenge is the integration of multimodal information. General-purpose multimodal pre-trained models, while adept at multimodal representation learning, struggle with fine-grained retrieval required for knowledge-intensive visual questions. Recent specialized pre-trained models mitigate the issue, but are computationally expensive. To bridge the gap, we propose a Unified Retrieval-Augmented VQA framework UniRVQA. UniRVQA adapts general multimodal pre-trained models for fine-grained knowledge-intensive tasks within a unified framework, enabling cross-task parametric knowledge sharing and the extension of existing multimodal representation learning capability. We further introduce a reflective-answering mechanism that allows the model to explicitly evaluate and refine its knowledge boundary. Additionally, we integrate late interaction into the retrieval-augmented generation joint training process to enhance fine-grained understanding of queries and documents. Our approach achieves competitive performance against state-of-the-art models, delivering a significant 4.7% improvement in answering accuracy, and brings an average 7.5% boost in base MLLMs’ VQA performance, all within a total training time of under 3 hours. The code is available at https://anonymous.4open.science/r/UniRVQA-D8C7.
Retrieval-augmented Generation, Multimodal Knowledge Reasoning, Knowledge-based Vision Question Answering
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Language modelsccs: Information systems Question answeringccs: Information systems Information systems applications
1. Introduction
Knowledge-based Vision Question Answering (KB-VQA) is a task of answering challenging image-grounded questions that require the integration of external knowledge beyond commonsense, such as encyclopedic documents from the web (Vrandečić and Krötzsch, 2014; Luo et al., 2021). One of the approaches to KB-VQA involves leveraging the implicit knowledge of large pre-trained models (Yang et al., 2022; Alayrac et al., 2022; Chen et al., 2023b; Ma et al., 2024). However, these large models are less flexible in updating with the latest knowledge and often fail to mitigate their inherent factual errors (Mallen et al., 2023; Min et al., 2023). Alternatively, Retrieval-Augmented Generation (RAG) methods have emerged as a more flexible and lightweight solution for the KB-VQA task. RAG-based systems achieve knowledge-intensive generation by relying on the supporting knowledge that retrieved from external sources, providing the generator with relevant context and reducing the need for large-scale tuning.
Figure 1. A comparison between the large-model-based framework, multi-stage RAG framework, and our proposed unified RAG framework (UniRVQA). The representative systems are listed in gray texts for exemplifications.
Existing work on RAG-based KB-VQA often employs sequential and independent models for the retriever and answer generator (Gui et al., 2021; Lin et al., 2022, 2023). Such modular separation inherently prevents the retriever and generator from benefiting each other’s training process and mutually sharing parametric knowledge. Recent efforts  (Lin and Byrne, 2022; Siriwardhana et al., 2023) have demonstrated promise in mitigating the issue by improving the retriever’s performance through shared optimization with the generator. However, cross-task interactions and effective parametric knowledge sharing are remain constrained. These approaches retain modules separated and operate unidirectional knowledge sharing: the retriever’s parametric knowledge is refined through the generator’s training signals, without providing reciprocal feedback or sharing parametric knowledge with the generator. On the other hand, knowledge retrieval and answer generation are highly related tasks, where both of them require the precise understanding and reasoning over external knowledge documents, visual questions context, and their interrelations. We argue that a more integrated framework, facilitating mutual parametric knowledge sharing, could better exploit the interdependence between tasks and lead to improved system performance.
The other fundamental challenge of KB-VQA lies in the multimodal representation learning. Earlier verbalization approaches often convert images into text (Lin et al., 2022; Gui et al., 2021; Gao et al., 2022), by image captioning and dense labeling for example. The vision-grounded task will then be reformulated as a standard textual question-answering problem. However, such transformation can result in the loss of complex visual details, thereby limiting the ability in context understanding and knowledge retrieval (Ma et al., 2024; Lin et al., 2023). With recent advancements in multimodal alignment techniques (Radford et al., 2021; Li et al., 2022), general pre-trained Multimodal Large Language Models (MLLMs), such as BLIP2 (Li et al., 2023) and InstructBLIP (Dai et al., 2023), have demonstrated their superior abilities in extracting multimodal representations and answering general visual questions. However, these pre-trained models often fail to achieve satisfactory retrieval and answering performance on knowledge-intensive bases because they are less capable of capturing fine-grained nuances. While efforts have been put into pre-training specialized models for effective knowledge-based multimodal retrieval and question answering (Chen et al., 2022; Yasunaga et al., 2022; Hu et al., 2022; Lin et al., 2023, 2024), it would be highly beneficial with promising potential to adapt existing general pre-trained multimodal models for RAG-based KB-VQA as they already memorized extensive general knowledge. Such adaptation requires much fewer data and computational resources compared to pre-trained specialized models, yet it is currently remains underexplored.
In response, this paper proposes the Unified Retrieval-Augmented Vision Question Answer (UniRVQA) framework. As demonstrated in Fig. 1-B, our framework enables a pre-trained MLLM to effectively handle all of tasks along the KB-VQA pipeline through shared parametric knowledge. Specifically, UniRVQA jointly optimizes the objectives of reflective-answering and retrieval-augmented generation. In the reflective answering branch, the model is trained to utilize its implicit knowledge repository for answering questions. A reflective-answering mechanism is proposed to enable the model to evaluate the correctness of its answers immediately after generation. During inference, this reflection mechanism determines whether external knowledge should be engaged. In the retrieval-augmented generation pathway, the process is framed as Bayesian joint probability prediction, which enables the simultaneous training of fine-grained knowledge retrieval and knowledge-dependent answer generation tasks. Additionally, we integrate late interaction mechanism (Khattab and Zaharia, 2020) into the joint training framework to further enhance fine-grained level information understanding. The experiment shows that the aforementioned tasks can complement each other, which results in improved performance in both retrieval and VQA. The main contributions of this paper are summarized as:
  • We propose a novel joint training framework that enables the general MLLM to be a unified framework, which can effectively handle both tasks along the KB-VQA pipeline, including knowledge-dependent visual question answering and fine-grained knowledge retrieval.
  • We introduce a novel reflective-answering mechanism that empowers the model to assess its knowledge boundary and adaptively perform RAG during inference.
  • Experiments on two public datasets show that our model outperforms state-of-the-art methods with significant improvements of 4.78% in answer accuracy and improves base MLLMs by an average of 7.54% in answering accuracy.
2. Relate Work
2.1. Knowledge-based VQA Systems
KB-VQA systems (Yang et al., 2022; Lin and Byrne, 2022; Lin et al., 2023; Yu et al., 2023; Salemi et al., 2023) solve complex knowledge-intensive visual questions. One of the approaches to KB-VQA is to leverage implicit knowledge from large language models, such as GPT-3 (Brown et al., 2020), using carefully crafted prompts. For example, KAT (Gui et al., 2021) and PICa (Yang et al., 2022) transform images into textual captions so that the visual context can be utilized as part of the prompts by GPT-3(Brown et al., 2020). However, these large models are less adaptable when updating with new knowledge and often fail to mitigate their inherent factual errors (Mallen et al., 2023; Min et al., 2023). Therefore, the retrieval-augmented generation (RAG) based paradigm has emerged as a more efficient and lightweight alternative for KB-VQA. RAG-based systems first retrieve relevant knowledge, which is then processed by the generator for answer generation. These external knowledge can either come from structured Knowledge Graphs (Li et al., 2020a; Guo et al., 2022; Speer et al., 2016) or unstructured documents such as documents on Wikipedia  (Vrandečić and Krötzsch, 2014; Lin et al., 2023).
Existing works on RAG-based KB-VQA often have independent retriever and generator models that are trained separately (Lin et al., 2022; Lin and Byrne, 2022; Lin et al., 2023; Hu et al., 2022). For example, REVIVE (Lin et al., 2022) adopts a pre-trained CLIP (Liu et al., 2023) to extract features for retrieval and adopt multiple FiD networks  (Izacard and Grave, 2021) as the backbone of the generator. However, the independence of the generator and retriever can lead to compromised performance as their tasks are closely related. To mitigate the issue, RA-VQA (Lin and Byrne, 2022) propose to first train the retrieval network, followed by joint training of both the generator and retriever. However, cross-task interactions and parametric knowledge sharing are still constraint with the separated modules. Given the interdependence across KB-VQA tasks, we suggest that it would be beneficial to have a more integrated framework as both tasks need a fine-grained understanding of knowledge and questions. Moreover, existing work (Lin and Byrne, 2022; Yang et al., 2022; Gao et al., 2022) often verbalize images by applying image-to-text transformation (Shi et al., 2017; Anderson et al., 2018; Li et al., 2020b), which often results in the loss of critical fine-grained visual information.
2.2. Pre-trained Multimodal Models in KB-VQA
Recent large vision-language pre-trained models  (Radford et al., 2021; Dai et al., 2023; Li et al., 2023) offer an effective solution to the above challenge brought by verbalization, with its superior capabilities in multimodal representation learning. These general MLLMs are usually pre-trained on large-scale datasets so that knowledge can be stored in model parameters. Therefore, MLLMs like BLIP2 (Li et al., 2023) have a superior visual understanding capability, by constructing a lightweight Querying Transformer between visual encoders (e.g. ViT-L/14 (Dosovitskiy et al., 2020)) and LLMs (e.g. Flan-T5 (Raffel et al., 2019)) in a wide range of downstream multimodal tasks, such as image-text retrieval and image-grounded question answering. However, they still struggle with document retrieval for KB-VQA and, therefore, fail to elevate themselves to a more satisfactory answering performance. This is because these general pre-trained models are not primarily trained to capture fine-grained nuances within external documents.
Figure 2. An overview of the proposed Unified Retrieval-Augmented Vision Question Answering framework (UniRVQA). The framework consists of two main pathways: (1) Part A and B perform late-interaction retrieval and retrieval-augmented generation, which together form the RAG path. (2) Part C outlines the reflective-answering mechanism, where the base model conducts self-answering and evaluates the correctness simultaneously.
To address this issue, recent efforts have focused on specialized pre-trained models to enhance their knowledge-intensive retrieval and answering capabilities. For instance, FLMR (Lin et al., 2023) designs a mapping layer to project visual embeddings to token-level language embeddings for downstream retrieval tasks. MuRAG (Chen et al., 2022) trains a cross-modal transformer to fuse the visual and textual embeddings. The model undergoes pre-training on a large-scale dataset that integrates images, text, and knowledge by applying the joint learning strategy. Similarly, RA-CM3 (Yasunaga et al., 2022) also injects knowledge bases during the pre-training process to align image-text-knowledge tuples so that the model can be equipped with knowledge retrieval capabilities. Despite the relatively promising performance these models have achieved, we argue that these pre-training methods are computationally expensive and less flexible to update with latest information. Given the rich parameterized knowledge of general MLLMs, it would be worthwhile to adapt them for their potential in KB-VQA, which is a non-trivial yet sparsely researched question.
3. Methodology
Before introducing the proposed framework, Section 3.1 provides a formal mathematical formulation of the KB-VQA task. To address the challenge of limited knowledge sharing between the retriever and generator, our method UniRVQA leverages a unified encoder-decoder structure built upon a base pre-trained MLLM. The encoder processes both textual and visual inputs into a unified semantic space. Section 3.3 details our novel self-reflective joint training approach, focusing on two key pathways: retrieval-augmented generation and reflective answering. Finally, we describe how answers are reasoned out, emphasizing how the reflective-answering mechanism facilitates adaptive RAG generation during inference.
3.1. Problem Formulation
We consider the general setting of KB-VQA for the framework design. Given a textual question $Q$ regarding an image $I$, the objective of KB-VQA is to generate an answer $\hat{a}$ based on retrieved relevant documents $\mathcal{D}=\{d_{i}\}^{k}_{i=1}$:
(1) $\hat{a}=\arg\max_{a,d_{i}\in\mathcal{D}}p_{\Phi}(a|Q,I,d_{i}),\$
where $\Phi$ denotes the parameters of the base model. The answer is based on the Bayesian joint probability of retrieval and generation:
(2) $p(a|Q,I,\mathcal{D}_{full})=\underbrace{p_{\phi}(\mathcal{D}|Q,I,\mathcal{D}_{full})}_{\text{Retrieval}}\cdot\underbrace{p_{\Phi}(a|Q,I,\mathcal{D})}_{\text{Generation}},$
where $\mathcal{D}_{full}$ represents the external knowledge base of size $N$ and $\phi$ denotes the parameters for retrieval models. In the UniRVQA setting, the number of retrieved documents $k\ll N$ and $\phi\subseteq\Phi$.
3.2. Unified Multimodal Embedding
With the pre-trained MLLM as our base model, we can first obtain the unified embeddings from both textual and visual input to construct a query embedding $\mathcal{\tilde{Q}}$:
(3) $\mathcal{\tilde{Q}}=[f_{mm}(Q),f_{mm}(I)]\in\mathbb{R}^{l_{Q}\times h}\$
where $h$ is the hidden size and $l_{Q}$ is the total length of the sequence by concatenating image tokens and text tokens embeddings. $f_{mm}$ is a part of the pre-trained MLLM that generates semantically-meaningful embeddings. In our framework, we adopt BLIP2 (Li et al., 2023) or InstructBLIP (Dai et al., 2023) as the base model, where $f_{mm}$ consists of a pre-trained Q-Former and a T5 encoder. With the same encoder, we can obtain the embedding of the document $d$ with a length $l_{d}$ in the external knowledge base:
(4) $\mathcal{\tilde{D}}=[f_{mm}(d))]\in\mathbb{R}^{l_{d}\times h}\$
3.3. Self-Reflective Joint Training
Instead of adopting the conventional framework where the retriever and generator are separate, we adopts a unified framework, which is trained under the proposed self-reflective joint training method to optimize for retrieval and answer generation simultaneously, allowing them to complement each other during training.
3.3.1. Late-Interaction Knowledge Retrieval
During the retrieval stage, the relevance $r$ between a particular query $\mathcal{\tilde{Q}}$ and a document $\mathcal{\tilde{D}}$ is assessed using a relevance score in a late-interaction manner, following ColBERT (Khattab and Zaharia, 2020). Late-interaction retrieval is known as a fine-grained and more efficient approach, where query and document representations are independently encoded before interacting (Khattab and Zaharia, 2020). We extend the ColBERT from a BERT-based retrieval model to any MLLM-based framework that can be jointly trained with the generation task:
(5) $r(\mathcal{\tilde{Q}},\mathcal{\tilde{D}})=\sum_{i=1}^{l_{q}}\max_{j=1}^{l_{d}}\mathcal{\tilde{Q}}_{i}\mathcal{\tilde{D}}_{j}^{T}$
The relevance score is calculated based on the token-level embeddings. For each token in the query, the document token with the highest relevance score will be identified, and these maximum scores will then be summed up to produce the overall relevance score. Our empirical study highlights that an additional step is needed to extend the use of ColBERT (Khattab and Zaharia, 2020), as the hidden size of the output from the MLLMs encoder is relatively large to conduct relevance calculation and indexing. To improve the efficiency of the retriever, we adopt a simple yet effective compression module that packs token embeddings into lower-dimensional latent spaces by using two multi-layer perceptron layers, connected by a ReLU activation function.
To elicit the retrieval ability of pre-trained MLLMs by learning the fine-grained relevance of the query and documents, we adopt an in-batch contrastive learning strategy, following (Karpukhin et al., 2020; Luo et al., 2021; Lin et al., 2023). Given each query $\mathcal{\tilde{Q}}$ in the batch, the ground-truth positive documents for other queries in the same batch will be considered as its negative samples, denoted as $D_{n}$. The contrastive learning retrieval loss will be formulated as:
(6) $\mathcal{L}_{R}=-\sum_{\mathcal{\tilde{Q}},\mathcal{\tilde{D}}^{+}}\log\frac{\exp(r(\mathcal{\tilde{Q}},\mathcal{\tilde{D^{+}}})}{\exp(r(\mathcal{\tilde{Q}},\mathcal{\tilde{D}}^{+})+\sum_{\mathcal{\tilde{D}}^{+}\in D_{n}}\exp(r(\mathcal{\tilde{Q}},\mathcal{\tilde{D}}^{+})}$
3.3.2. Retrieval-Augmented Generation
After the late-interaction relevance calculation, the embeddings of the positive document will be concatenated with the query embeddings and fed to the language model for answer generation. The concatenated retrieval-augmented embedding will be denoted as $y$ and the answer generation will be trained by the casual language modelling loss:
(7) $\mathcal{L}_{RAG}=-\sum_{i=1}^{l_{a}}y_{i}\log p(\hat{y}_{i}|y_{<i}),$
where $l_{a}$ is the length of ground truth answer $\mathcal{A}$. For each open-ended question, there will be a set of human responses $S$. The target answer will be randomly selected from the set. Our empirical studies show that random selection enhances the model’s robustness to document noise compared to using the answer provided in the document. Based on Eq. 2, the two loss terms can be directly summed together as the log-joint probability of retrieval and generation:
(8) $\mathcal{L}_{RAG\_joint}=\mathcal{L}_{R}+\mathcal{L}_{RAG}$
Algorithm 1 Pseudo Code of UniRVQA Reflective Answering Training Process
1:MLLM (Encoder $\mathcal{E}$, $\mathcal{D}$, Parameters $\Phi$); Question-image pairs $q$; batch $\mathcal{B}$ with size $n$; Target Answer $\mathcal{A}$; Learning rate $\alpha$
2:$\mathcal{L}_{SR}=\mathcal{L}_{gen}+\mathcal{L}_{reflect}$
3:for t=1,2,…,T do
4:     for each image-question pair ($q_{i}$) in $\mathcal{B}_{t}^{n}$ do
5:         Predict answer: $\hat{a_{i}}=\mathcal{D}(\mathcal{E}(q_{i}))$
6:         Calculate the loss for the self-generation:
7:         $\mathcal{L}_{SR}=\mathcal{L}_{gen}=\sum_{i}^{n}\textsc{CrossEntropyLoss}(a_{i},\mathcal{A}_{i})$
8:         while $t\geq s_{join}$ $\triangleright$ Late join
9:              Construct self-reflection label $r_{i}$:
[…]
12:              end if
13:              Predict self-reflection result: $\hat{r_{i}}=\mathcal{D}(\mathcal{E}(q_{i}\oplus\hat{a_{i}}))$
14:              Calculate the loss for the self-reflection:
15:              $\mathcal{L}_{reflect}=\sum_{i}^{n}\textsc{CrossEntropyLoss}(\hat{r_{i}},r_{i})$
16:              Update: $\mathcal{L}_{SR}=\mathcal{L}_{gen}+\mathcal{L}_{reflect}$
17:         end while
18:         Update encoder-decoder parameters $\Phi$
19:     end for
20:end for
3.3.3. Reflective Answering
Through our preliminary experiments, we found that the standard training process, as described in Eq. 7, allows the model to retrieve from documents but also encourages excessive reliance on external knowledge, even when the information is less relevant. This overreliance may result in the model inevitably embedding noise into its parametric knowledge during training. On the other hand, pre-trained MLLMs already possess the required knowledge to directly answer some easier knowledge-intensive questions. Therefore, we propose a novel on-the-fly Reflective-answering mechanism. which trains the generator without external documents while simultaneously generating a self-reflection label. The self-reflection label indicates whether the model considers its answer to be correct based on the context of the question:
(9) $p_{\Phi}(\text{"}correct"|Q,I)=p_{\Phi}(\hat{a}|Q,I)\cdot p_{\Phi}(\text{"}correct"|Q,I,\hat{a})$
Every time when the answer $\hat{a}$ is generated, the self-reflection label can be generated immediately by comparing $\hat{a}$ with the target answer $\mathcal{A}$. The generated answer will then be concatenated with the query to serve as the context for the binary label prediction – “correct” or “incorrect”. In this setting, the binary classification task is framed as the next-token prediction, where another casual language modelling loss following the Eq. 8 will be calculated on self-answer and self-reflection generation, denoted as $\mathcal{L}_{SR}$. The joint loss will then be updated with $\mathcal{L}_{SR}$ in Eq. 10. Formally, the reflective answering path is described in Alg. 1.
(10) $\mathcal{L}_{Joint}=\mathcal{L}_{R}+\mathcal{L}_{RAG}+\mathcal{L}_{SR}$
[…]
Train 8,112 33,212
Validation 912 2,000
Test 5,046 73,620
[…]
# doc for training 111,411 100,000
# doc for testing 166,389 100,000
Table 2. Model performance comparison on OK-VQA. The best performance of our model is highlighted in bold font, and the rows of our models’ main results are gray. The best performance in literature is underlined. $K$ is the amount of knowledge retrieved in the generation process. Know. Source represents external knowledge source.
No. Model Base Models K Know. Source EM(%) VQA(%)
Classic KB-VQA Systems
1 KAT-T5 T5-Large 40 Wikipedia - 44.25
2 TRiG T5-Large 100 Wikipedia 54.73 50.50
3 MAVEx - - Wikipeida - 39.20
4 RA-VQA T5-Large 5 GoogleSearch 55.77 51.22
5 BLIP (zero-shot) BLIP - - 36.99 34.46
6 BLIP (fine-tuned) BLIP - - 48.89 45.74
[…]
Systems with Large Models (¿15B parameters)
8 PICa GPT-3 (175B) - - - 48.00
9 Prophet GPT-3 (175B) - - - 61.11
10 REVIVE GPT-3 (175B) 40 Wikipedia - 58.00
11 PALI PALI (15B) - - - 56.50
12 Flamingo Flamingo (80B) - - - 57.80
13 PaLM-E PaLM-E (526B) - - - 66.10
Base Models without Knowledge Retrieval
[…]
Our Proposed Models ($\sim$3B parameters)
[…]
3.4. Inference and Answer Select
At the inference stage, all documents will first be indexed using PLAID (Santhanam et al., 2022) for accelerated late-interaction retrieval. The self-reflection mechanism not only enables the model to evaluate its knowledge boundaries but also facilitates adaptive retrieval-augmented generation (RAG) during the UniRVQA inference stage. At this stage, the model first answers the question without referencing external documents and immediately generate self-reflection prediction to assess the correctness of its response. If the model deems its answer incorrect, the RAG process is triggered, allowing it to retrieve documents to facilitate answering. Otherwise, the self-answering result will be kept as the final answer. During the retrieval-augmented generation stage, multiple documents will be selected and generate multiple answers accordingly. Based on Eq. 1 and  2, the answer with the highest joint probability of retrieval and generation will be selected.
4. Experiments and Results
4.1. Experiments Setup
Datasets. We mainly evaluate the proposed method on the OK-VQA dataset (Marino et al., 2019) and conduct complementary experiments using the InfoSeek dataset (Chen et al., 2023a) to demonstrate the model’s generalizability. InfoSeek is regarded as more knowledge-dependent than earlier KB-VQA datasets, as its questions are often unanswerable without external knowledge support. Here are the details about two datasets:
(1) OK-VQA (Marino et al., 2019) dataset contains over 10k questions on MSCOCO (Lin et al., 2014) images which require external knowledge to answer. For the external knowledge base we adopt Google Search Corpus (Luo et al., 2021), which is a textual corpus, containing 166,389 passages from Google, covering all the knowledge necessary for answering questions in OK-VQA. We use the original splits of training and testing sets to ensure comparability, and use 10% of the training set for validation.
(2) InfoSeek (Chen et al., 2023a) is a newly proposed large-scale KB-VQA dataset introduced in 2023, built on the OVEN image dataset (Hu et al., 2023). Following the original paper we use Wikipedia (Vrandečić and Krötzsch, 2014) as the external knowledge base. Compared to OK-VQA, InfoSeek is more challenging as it encompasses a greater number of questions that necessitate expertise knowledge for accurate responses. Given the large size (over 1 million image-question pairs) of InfoSeek and Wikipedia corpus, we conduct our experiments on a down-sampled subset following existing works (Lin et al., 2024). Specifically, to reduce the large number of duplicate samples while maintaining its diversity, we stratified the training set by the combination of entity ID and question, and randomly sampled $\sqrt{n_{i}+1}+1$ entries from $n_{i}$ entries in each group. This sampling operation reduced the training set of Infoseek from  934k to  33k. To ensure comparability with prior works (Lin et al., 2024), we used the original validation set as the test set and randomly sample 2,000 entries from the original training set as the validation set. We also randomly sampled 100k documents from the 6M wikipedia corpus while making sure of the presence of relevant documents. Details of the data splits statistics for both datasets can be found in Table 1.
Implementation Details. We select BLIP2-Flan-T5-XL (Li et al., 2023) and InstructBLIP-Flan-T5-XL (Dai et al., 2023) as the MLLM base models to build our proposed framework. We use 1 Nvidia A100 with 80GB VRM for all experiments. We use DoRA (Liu et al., 2024) to fine-tune UniRVQA. We choose a batch size of 20 and the AdamW optimizer (Loshchilov and Hutter, 2017) with the learning rate set as 2e10-4. The scheduler modulates the learning rate throughout the training process, starting with a warmup period of 100 steps before gradually reducing the learning rate following a cosine schedule. We evaluated the models on the validation every 200 steps and the best performing checkpoints were found to be step 2800 (OK-VQA) and step 3400 (InfoSeek). The documents were truncated to a maximum length of 256 tokens. The beam size was set at 3 for answer generation. To ensure comparability and avoid randomness bias, we report our main results as the average from 3 different random seed settings.
Computation Costs. Given the best-performing checkpoints on two datasets, We report the required total training hours on 1 Nvidia A100 (80G) are on average less than 3 GPU hours. Compared to the comparable models also trained with 1 A100 and takes more than 20 GPU hours for the training progress (Lin et al., 2024, 2023), our method are less computationally expensive while achieving competitive effectiveness.
Evaluation. We present the metrics used to assess answer generation and knowledge retrieval performance:
(1)Exact Match (EM): We evaluate the exact matching between the generated answer and the answer set $S$, where #$\#s(\hat{a})$ is the occurrences of $\hat{a}$ in $S$:
(11) $\mathrm{EM}(\hat{a},S)=\mathrm{min}(\#s(\hat{a}),1)$
(2)VQAScore: on OK-VQA dataset, we use an additional official VQA Score (Marino et al., 2019). This score makes the model partially rewarded if it generates a less popular answer among human responses:
(12) $\mathrm{VQAScore}(\hat{a},S)=\mathrm{min}(\#s(\hat{a})/3,1)$
(3)Pseudo Relevance Recall (PRR@K): Following previous work  (Lin and Byrne, 2022; Lin et al., 2023), we adopt pseudo-relevance labels and evaluate the retrieval performance by counting the number of questions that successfully retrieve documents with correct answers contained in top-k retrieval results.
Baseline Models. We compare our proposed framework with the latest KB-VQA systems in answer generation performance. Among them, the first group of systems smaller model with less than 3B parameters, including:
  • KAT (Gui et al., 2021) integrates implicit and explicit knowledge into a transformer-based KB-VQA system to jointly reason over both knowledge sources.
  • TRiG (Gao et al., 2022) transforms all visual information into language space on three levels, including image-level captioning, object-level dense labeling and text OCR.
  • MAVEx (Wu et al., 2022) enhances KB-VQA performance by validating generated answers through retrieving multimodal external evidence and providing multimodal explanations to support its reasoning process.
  • RA-VQA (Lin and Byrne, 2022) also transforms visual content into textual space and proposes a joint training framework for retrieval and answer generation.
  • BLIP (Li et al., 2022) is one of the leading pre-trained multimodal models, designed to perform vision-language tasks such as VQA by leveraging unified image-text semantic space.
  • FLMR (Lin et al., 2023) proposes a pre-trained framework that focuses on fine-grained alignment between multimodal inputs, enabling high-accuracy retrieval and reasoning for KB-VQA.
The second group includes larger systems that are built with large pre-trained models such as GPT-3 (175B) (Brown et al., 2020) and PaLM-E (526B):
  • PICa (Yang et al., 2022) also conducts image-text transformation and prompts the GPT-3 for implicit knowledge.
  • Prophet (Yu et al., 2023) proposes heuristics-enhanced prompting to utilize the implicit knowledge of frozen LLMs.
  • REVIVE (Lin et al., 2022) exploits object-centric regional information together with image-question query to retrieve knowledge from GPT-3 and Wikidata (Vrandečić and Krötzsch, 2014).
  • PALI (Chen et al., 2023b), Flamingo (Alayrac et al., 2022) and PaLM-E (Driess et al., 2023) represent state-of-the-art large pre-trained Multimodal models that achieve strong performance on KB-VQA datasets.
Additionally, we will compare our retrieval performance against strong retrieval-focused models designed specifically for VQA task, mainly including Dense Passage Retrieval (DPR) (Karpukhin et al., 2020), FLMR (Lin et al., 2023) and preFLMR (Lin et al., 2024). FLMR enhances retrieval by leveraging ColBERT (Khattab and Zaharia, 2020) for fine-grained alignment, while preFLMR extends FLMR’s approach by pretraining on a substantially larger retrieval corpus to improve performance further.
Table 3. VQA performance comparison on original InfoSeek validation split with K=5. Un-Q and Un-E stand for two types of test questions – unseen questions and unseen entities.
Accuracy (%)
Model Base Models Un-Q Un-E All
Standard fine-tuned on the dataset
PaLM(Q-only) PaLM 5.5 4.2 4.8
[…]
PALI-17B PALI 24.2 16.7 19.7
Fine-tuned with knowledge
CLIP + PaLM PaLM(540B) 22.7 18.5 20.4
CLIP + FiD - 23.3 19.1 20.9
UniRVQA InstructBLIP 24.01 20.40 22.06
[…]
Figure 3. Retrieval performance variation with respect to the number of retrieved knowledge evaluated on OK-VQA.The DPR results shown are for baseline reference.
4.2. Main Results and Analysis
According to the experiment results, our key observations are:
  • UniRVQA significantly and consistently improves the VQA performance of base MLLMs, achieving an average accuracy gain of 7.42% and 7.66% on two datasets, demonstrating its ability to extend the effectiveness of existing MLLMs in knowledge-intensive tasks.
  • Our proposed model achieves the state-of-art performance in both answer generation and knowledge retrieval. It achieves the highest EM of 66.79% on the OK-VQA dataset, with a notable 4.78% improvement over the best model in literature. It also delivers the best knowledge retrieval results across datasets, with an average improvement over 3%.
  • Compared to other high-performing very large models, such as PaLM-E (526B)(Driess et al., 2023) and PaLI (15B) (Chen et al., 2023b), UniRVQA achieves competitive performance with a compact size (3B) and a training time of under 3 hours, demonstrating its high efficiency without sacrificing performance.
4.2.1. VQA Performance
The overall accuracy performance comparison between our models and the baseline models on the OK-VQA dataset is shown in Table 2. First, our proposed framework implemented with various base models (InstructBLIP (Dai et al., 2023) and BLIP2 (Li et al., 2023)) achieves the top-tier performance. Specifically, UniRVQA (InstructBLIP) delivers the best EM accuracy, with a 4.78% improvement over the previous best model (FLMR (Lin et al., 2023)). Additionally, it achieves a highly competitive VQAScore accuracy of 61.57%. Meanwhile, UniRVQA (BLIP2) secures second place in EM (64.21%) and maintains strong VQA accuracy at 60.90%. UniRVQA also demonstrates exceptional performance in complementary experiments on InfoSeek (Table 3). It achieves the highest answering accuracy of 24.01% on Unseen Questions and 20.40% on Unseen Entities.
Boosting base MLLMs performance. While extremely large models with more than 15B parameters(Table 2, lines 9–13) are initially advantageous in VQA performance and benefit from their capacity to store extensive knowledge through large-scale training, UniRVQA effectively closes the performance gap between them and smaller base models (3B) (Table 2, lines 14–17). By integrating UniRVQA, the base MLLMs achieve leading performance levels. Compared to their standard fine-tuned counterparts, the variants of UniRVQA deliver an average improvement of 7.42% (Table 2, lines 18–19 vs. lines 15, 17). The similar observation can be concluded from the InfoSeek dataset (Table 3), where UniRVQA raises the base model’s overall accuracy from 14.5% to 22.1%, demonstrating an impressive 7.56% gain. The outcomes highlight UniRVQA can greatly unlock the potential of general MLLMs in addressing knowledge-intensive VQA tasks. Its adaptable framework not only maximizes the utility of existing models but also paves the way for leveraging future advancements in MLLMs. Additionally, UniRVQA proves especially advantageous in scenarios with limited access to extremely large models or constrained computational resources.
Lastly, we would like to point out the performance on CLIP combined with FiD (Table 3), as referred from the original paper (Chen et al., 2023a), achieves the second-best result using a special setting of retrieving 100 documents per question, which is far exceeds 5 documents used by ours and most of baseline methods. Obtaining better performance under less favorable settings again highlights the effectiveness of our approach. Furthermore, our model achieves competitive results with remarkable efficiency, requiring only 3 GPU hours for 3,000 training steps, in contrast to other models that typically demand over 24 GPU hours (Lin et al., 2024, 2023; Driess et al., 2023; Chen et al., 2023b).
4.2.2. Retrieval Performance
Besides question-answering performance, we evaluate the model’s retrieval performance to understand the system’s ability in leveraging external resources. As shown in Table 4, UniRVQA (InstructBLIP) achieves state-of-the-art retrieval performance, with PRR@5 scores of 90.47% on OK-VQA and 68.51% on InfoSeek. While UniRVQA maintains its leading position on OK-VQA, we observe relatively small performance gaps between models on this simpler dataset, with UniRVQA surpassing the previous best model by only around 1% across retrieval levels. However, on the more challenging InfoSeek dataset, our model demonstrates significantly superior robustness, achieving an 8.91% improvement over existing methods, which typically struggle with the performance drops. This underscores UniRVQA’s strong capability to tackle complex, knowledge-intensive scenarios.
Additionally, we analyze how retrieval performance evolves with the number of retrieved knowledge passages. Figure 4 shows that UniRVQA (InstructBLIP and BLIP2) achieves high recall earlier in the retrieval process, reaching approximately 80% recall with only the top-2 passages, whereas the baseline requires the top-5 passages for comparable performance. Beyond this, UniRVQA improves sharply by 14% as K increases from 1 to 5, after which improvements taper off. This demonstrates that the model efficiently retrieves relevant documents, reducing the computational burden of excessive retrievals. Furthermore, UniRVQA’s superior retrieval performance highlights that, with the unified framework, reasoning skills acquired during answer generation can positively impact retrieval. This synergy is further explored in the ablation study.
Table 4. Retrieval performance comparison on InfoSeek and OK-VQA. We report only PRR@5 on InfoSeek to follow the previous work. The best performance of our model is highlighted in bold font.
InfoSeek OK-VQA
Model PRR@5 PRR@5 PRR@10
DPR 44.88 82.90 89.95
FLMR 46.42 89.32 94.00
PreFLMR 59.60 - -
UniRVQA(BLIP2) 63.65 88.35 92.69
UniRVQA(InstructBLIP) 68.51 90.47 94.53
4.3. Ablation Study
Additionally, we analyze the effectiveness of main components in our proposed framework to answer the following research questions. All experiments are conducted on the OK-VQA dataset.
Figure 4. Ablation study on the self-reflection mechanism. Two groups of models in each graph are based on InstructBLIP and BLIP2 respectively.
RQ1: How does joint training on a unified framework work? To investigate how does joint training on a unified framework affects performance, we construct two variants of UniRVQA by separating the training process of retriever and answer generator in different ways. We keep the late interaction and reflective-answering mechanism in all variants. The results are summarized in Table 5. Specifically, $\mathrm{UniRVQA^{-}}$ still uses a unified framework which will first be trained on the retrieval task, followed by training on the answer generation task. $\mathrm{UniRVQA^{--}}$ employs the more traditional setting of two separate models as the retriever and the generator. By training the model separately as a retriever and a generator, $\mathrm{UniRVQA^{--}}$ performs relatively well in all metrics, although slightly worse than the full model. This indicates that base MLLMs have the potential to handle both knowledge-intensive tasks separately. In the proposed unified framework, the retriever and answer generator further share the same network, reducing model size significantly and allowing both tasks to complement each other, resulting in improved performance.
When inspecting the performance from $\mathrm{UniRVQA^{-}}$, the results show that simply unifying the framework without proper training strategy design leads to a significant drop in performance. This suggests that naive multistage training may hinder the model’s ability to share parametric knowledge effectively. The better performance on UniRVQA verifies our assumption that by sharing parametric knowledge across tasks, the model can better leverage capabilities learned from one task to improve the performance on the other.
[…]
Model PRR@5 VQA(%) EM(%)
$\mathrm{UniRVQA^{--}}$ 86.45 57.97 60.01
$\mathrm{UniRVQA^{-}}$ 76.00 49.13 51.38
UniRVQA 90.47 61.57 66.79
RQ2: How can the Reflective-Answering enhance the system performance? To investigate the impact of reflective-answering mechanism, we conduct ablation studies where the mechanism is disabled during the training and inference stage respectively. The results, presented in Fig. 4, show that removing reflective-answering leads to a significant drop in performance compared to the complete UniRVQA model. Specifically, removing self-reflection during training results in the most substantial decrease, with a reduction of over 6.91% in VQA score and more than 7.5% in EM for both base models. This highlights the critical role of self-reflection in improving model effectiveness during training. In addition, turning off self-reflection during inference hinders performance by approximately 1.3%, further demonstrating that reflective-answering is crucial not only for learning but also for making more accurate predictions. Therefore, encouraging the model to wisely rely on its implicit knowledge during inference yields accuracy gains ranging from 1.22% to 3.03%. We attribute the effectiveness of our proposed reflective-answering mechanism to its ability to alleviate the model’s dependence on external knowledge. This mechanism reminds the model to leverage its own implicit knowledge, which has been empirically proven to be useful (Yang et al., 2022; Gui et al., 2021). With reflective-answering, the model is more discerning in its use of external information, avoiding reliance on irrelevant data when it already possesses sufficient implicit knowledge to answer accurately.
RQ3: How does the external knowledge support the proposed system? Here, we further investigate the contributions of external knowledge in our experiments firstly by removing them during the inference, as in Figure 4. We conclude that removing the external knowledge support will directly lead to on average 4.1% reductions in answer accuracy, which confirm the necessity of external knowledge when answering knowledge-intensive questions. Compared the standard fine-tuning method that provide the model with final answer as training signals (Table 2, line 15 and 17), UniRVQA framework brings about 2% improvement in answering accuracy. Specifically, UniRVQA (InstructBLIP) improves the fine-tuned counterparts from 60.47% to 62.64% in EM and from 55.50% to 57.38% in VQAScore. Such improvement indicate that incorporating the external knowledge during the training process appropriately can also be beneficial to enhance the reasoning ability of answer generator. The theoretical mechanism is also worth to be investigated in the future research.
4.4. Case Studies
[…]
The right column features two examples from InfoSeek. The top-right example asks for the specific location where the facility is standing, a question that would be very hard to answer without particular external knowledge. This also highlights the complexity of questions in InfoSeek. UniRVQA accurately retrieves an encyclopedic document describing the ferris wheel installed on the “Place de la Concorde” in Paris, showcasing the model’s ability to effectively identify and use fine-grained information to answer challenging questions. We also note that although the base model was not able to provide the precise location, it could still identify the city in the image, which indicates that the base MLLM contains some fundamental knowledge that can be potentially leveraged. The bottom-right failure case is challenging. The model struggled with identifying the species of the black bird in the image, however still managed to retrieve generally relevant information about “blackbird”. We emphasize that a potential area for improvement is the fine-grained entity retrieval, particularly in distinguishing visually similar entities.
5. Conclusion and future work
In this paper, we propose a Unified Retrieval-Augmented Vision Question Answer framework (UniRVQA), which can effectively adapts general-purpose MLLMs for both fine-grained retrieval and knowledge-intensive answer generation tasks, through a self-reflective joint training framework and incorporating a reflective-answering mechanism to optimize implicit and explicit knowledge utilization. Extensive experiments demonstrate that UniRVQA can elevate underperforming base models to leading positions. UniRVQA also achieve competitive answering and retrieval performance, compared to state-of-the-art models, all while maintaining a smaller model size and being training efficient—making KB-VQA research more accessible. The ablation study shows that the unified framework can enhance the performance by sharing parametric capabilities to complement tasks along RAG.
In the future, we aim to enhance system performance by focusing on more accurate entity retrieval and recognition. One key direction will involve refining the model’s ability to identify and extract relevant entities from large knowledge bases, as this plays a critical role in generating contextually precise answers. The other direction lies in identifying the mechanism that external knowledge works to help with the life cycle of Knowledge-intensive VQA systems.